# Tunable non-Markovian dynamics in a collision model: an application to coherent transport

### v2

@@ Abstract
We propose a collision model to investigate the information dynamics of a system coupled to an environment with varying degrees of non-Markovianity. We control the degree of non-Markovianity by applying a depolarising channel to a fixed and rigid reservoir of qubits. We characterise the effect of the depolarising channel and apply the model to study the coherent transport of an excitation on a chain of three interacting qubits. We show how the system-environment coupling strength and the degree of non-Markovianity affect the process. Interestingly, in some cases a Markovian environment is preferable to enhance the coherent transport of the excitation.
**footnotetext: These authors contributed equally to this work.
@@ I Introduction
In this paper, we take an information-theoretic approach to investigate coherent transport in a simplified scenario. We propose a collision model [27, 28] to study the non-Markovian information dynamics of a chain of qubits interacting with the environment. Differently from other models in the literature [29, 30, 31, 32, 33, 34, 35, 36], the system interacts locally with the same reservoir qubits and we control the degree of non-Markovianity of the environment by applying a depolarising channel to the reservoir. By changing the intensity of the depolarising effect, we can control the amount of information irretrievably lost by the environment.
We first quantify the degree of non-Markovianity that the depolarising channel induces in the system. Then, we use the collision model to study the coherent transport of an excitation along a chain of qubits coupled to a fixed and rigid reservoir of interacting qubits. By changing the degree of non-Markovianity of the reservoir, we can model different types of environments. For example, a Markovian environment could represent a thermal bath surrounding the chain, while a non-Markovian one could model other degrees of freedom of the chain that are coupled to the ones responsible for the coherent transport.
We show that the transfer of information to the environment plays a crucial role in the coherence of the excitonic transport. The process displays different behaviours depending on the strength of the system-environment coupling. When the system-environment coupling is strong, non-Markovianity dramatically improves the coherence of the transport. Interestingly, when the coupling is weak, a Markovian environment is preferable to enhance the amount of coherence and the coherence time of the excitation.
Our results provide some information-theoretic insights into how non-Markovian environments affect the coherence of the excitonic transport on a chain. Moreover, our model demonstrates how to introduce a control on the degree of non-Markovianity in collision models with a fixed environment and might be applied in other fields where energy transport phenomena are relevant.
@@ II The model
We model the environment as a reservoir of $N$ interacting “reservoir qubits” $Q_{m}$, $m=1,...,N$. Their interaction is described by a Heisenberg Hamiltonian:
@@
with coupling constant $J^{m}_{res}\in\mathbb{R}\ \forall m=1,...,N$ and $\hat{\mathcal{X}}_{m}$, $\hat{\mathcal{Y}}_{m}$, $\hat{\mathcal{Z}}_{m}$ the Pauli operators of the m-th reservoir qubit. Since we are interested in how the information dynamics of the system is affected by a loss of information, we initialise the reservoir qubits in the maximally mixed state
@@
where $\mathbb{I}$ is the identity operator, $\hat{\mathcal{S}}_{m}$ the swap operator between the m-th system and reservoir qubit, and $\eta\in[0,\pi/2]$ a parameter that regulates the strength of the interaction. Since $\hat{P}_{m}(\eta)$ can be written as
@@
all the interactions in the model up to now are fundamentally of the same type.
Non-Markovianity.
Since all the interactions considered are unitary and the same reservoir qubits keep interacting with the chain, the model is intrinsically non-Markovian. Unlike other non-Markovian collision models where the reservoir qubits that have interacted with the system are discarded at some point [32, 33, 34, 35, 36], we control the degree of non-Markovianity of the model by applying a quantum depolarising channel to the reservoir qubits after each collision. This allows us to vary the degree of non-Markovianity in a fixed and rigid environment.
If $\rho_{CR}$ is the state of chain+reservoir, the depolarising channel for a single reservoir qubit can be expressed as:
@@
The parameter $\Omega$ controls the amount of information the environment loses. When $\Omega=0$, $\Delta_{0}^{m}(\rho_{CR})$ leaves $\rho_{CR}$ unchanged so that the global evolution of chain+environment is unitary. When $\Omega=1$, instead, $\Delta_{1}^{m}(\rho_{CR})$ re-initialises the m-th reservoir qubit to the maximally mixed state and, if applied to all the reservoir qubits, makes the environment Markovian. It is reasonable to assume that changing $\Omega\in[0,1]$ allows the exploration of different degrees of non-Markovianity of the environment. In Sec. III.1 and IV.1, we will show how the value of $\Omega$ directly relates to a measure of non-Markovianity in the model considered here. Since the degree of non-Markovianity is typically correlated with the environment’s temperature, we call $\Omega$ the effective temperature of the environment.
@@
Figure 1: Four steps of the protocol with three chain qubits and three reservoir qubits. a) exchange phase: each system qubit interacts with its nearest reservoir qubit via a partial swap; b) depolarisation phase: the depolarisation channel is applied to the reservoir qubits; c) transfer via reservoir: unitary evolution of the reservoir qubits with $\hat{H}_{res}$; d) transfer via chain: unitary evolution of the system qubits with $\hat{H}_{chain}$.
@@
    The system qubits interact with the nearest reservoir qubits via a partial swap (Eq. (4)) with strength $\eta$ (Fig. 1a). We will call this step the “exchange phase”.
@@
    The depolarisation channel is applied to the reservoir qubits (Fig. 1b). This step controls the degree of non-Markovianity of the environment. We will call this step the “depolarisation phase”.
@@
III Methods
In this section, we start by discussing the measure of non-Markovianity we use to characterise the effect of the depolarising channel. Then, we show how our model can be applied to study the coherent transport of an excitation along the chain.
III.1 Quantifying the degree of non-Markovianity
The measure of non-Markovianity we use is the one proposed in [38]. Given two initial states of the system $\rho_{1}(0)$ and $\rho_{2}(0)$, the degree of non-Markovianity under the action of a dynamical map is measured as [38]:
@@ III.1 Quantifying the degree of non-Markovianity
$\mathcal{N}=\underset{\{\rho_{1,0},\rho_{2,0}\}}{\text{max}}\sum_{n\in\Sigma_{+}}\left(D\left(\rho_{1,n},\rho_{2,n}\right)-D\left(\rho_{1,n-1},\rho_{2,n-1}\right)\right),$ (9)
where the summation is taken on all values of $n$ such that $D\left(\rho_{1,n},\rho_{2,n}\right)-D\left(\rho_{1,n-1},\rho_{2,n-1}\right)>0$, $\rho_{1,n}$ is the state of the system at the end of the $n$-th step obtained by tracing out the environment, and the maximisation is performed over all pairs of initial states $\{\rho_{1,0},\rho_{2,0}\}$.
@@
$\rho_{k,0}=\frac{1}{2}\begin{pmatrix}1+r_{k}\,\text{cos}\left(\theta_{k}\right)&r_{k}\,\text{sin}\left(\theta_{k}\right)e^{-i\Phi_{k}}\\ r_{k}\,\text{sin}\left(\theta_{k}\right)e^{i\Phi_{k}}&1-r_{k}\,\text{cos}\left(\theta_{k}\right)&\end{pmatrix},$ (10)
with $r_{k}\in\left[0,1\right]$, $\theta_{k}\in\left[0,\pi\right]$, and $\phi_{k}\in\left[0,2\pi\right]$, for $k=1,2$.
@@
The maximisation procedure becomes too complicated with a higher number of qubits. In this case, we use some of the insights on the states maximising Eq. (9) drawn from the 1x1 case to choose a suitable pair of initial states.
III.2 Application to coherent transport
@@
We apply the model described above to study the coherence of the transport of an excitation along the chain. We represent an excitation localised on the first qubit as
@@
in the Hilbert space $\mathscr{H}_{S_{1}}\otimes\dots\otimes\mathscr{H}_{S_{N}}$. Applying the protocol described above, the excitation spreads through the chain and the environment. After repeating the protocol $k$ times, we trace out the environment and extract the coherence between the first and last qubit of the chain $\rho_{C}^{1,N}$, that is, the coefficient of the term |$\ket{100\dots 0}\bra{00\dots 01}$ | of the density matrix. We take $|\rho_{C}^{1,N}|$ as an indicator of the coherent transport efficiency across the chain.
@@
For simplicity, we consider a model with three chain qubits. We implemented the protocol as a quantum circuit in Qiskit [39] as follows. The exchange phase consists of a sequence of local unitaries $\hat{P}(\eta)$ as in Eq. (4) for each of the system qubits $S_{i}$; the depolarisation phase is implemented as a depolarising noise model on the identity gate, which applies the quantum channel of Eq. (6); the transfer phases are implemented as unitaries on system and reservoir with a time step $\Delta t=0.01$. We choose $J^{n}_{chain}=J_{chain}=10,\,\forall n$ and $J^{m}_{res}=J_{res}=1,\,\forall m$. This choice represents a scenario where the chain is the most efficient medium for the propagation of the excitation.
We first apply the protocol two times and study how $|\rho_{C}^{1,3}|$ varies for different values of $\eta$ and $\Omega$. This leaves enough time for the excitation to reach the last qubit of the chain, passing through the intermediate qubit, and is sufficient to obtain some general insights on how the environment affects coherent transport. Then, we select some values of $\eta$ and $\Omega$ and repeat the protocol with more iterations. The results show how $|\rho_{C}^{1,3}|$ varies over time.
IV Results
Here, we present the results obtained from the simulations of the collision model with the methods described in Sec. III.1 and III.2.
IV.1 Non-Markovianity
Figure 2: Degree of non-Markovianity $\mathcal{N}$ as a function of $\Omega$ for two different models. $\mathcal{N}$ is calculated over 20 time steps with $\eta=\pi/2$. The blue circles correspond to values of $\mathcal{N}$ calculated with 1000 randomly generated pairs of initial system states in the 1x1 model. The red circles are the values of $\mathcal{N}$ obtained with the initial states |$\ket{0}$ and |$\ket{1}$. The green squares correspond to the values of $\mathcal{N}$ for the 3x3 model with initial states |$\ket{000}$ and |$\ket{111}$.
In Fig. 2, we plot the degree of non-Markovianity $\mathcal{N}$ induced by our protocol in two models with different numbers of qubits. The circles refer to a model with one system qubit and one reservoir qubit (1x1), and the squares to the model with three system qubits and three reservoir qubits (3x3). We have calculated $\mathcal{N}$ over 20 temporal steps with $\eta=\pi/2$ for different values of $\Omega$.
The blue circles correspond to values of $\mathcal{N}$ for 1000 randomly generated pairs of initial states of the system in the 1x1 model. The red circles correspond to the initial states |$\ket{0}$ and |$\ket{1}$. The graph provides strong numerical evidence that |$\ket{0}$ and |$\ket{1}$ maximise $\mathcal{N}$ in Eq. (9). During the course of the simulations, we noticed that any two antipodal states on the Bloch sphere would maximise $\mathcal{N}$. This observation leads us to consider |$\ket{000}$ and |$\ket{111}$ as good candidates to maximise $\mathcal{N}$ in the 3x3 model. The green squares correspond to values of $\mathcal{N}$ for this pair of initial states and show a behaviour consistent with the 1x1 case. Due to the computational demand of simulating the 3x3 model with many different initial states, we consider only the pair {|$\left\{\ket{000},\ket{111}\right\}$. The green squares can then be interpreted as a lower bound on the non-Markovianity of the 3x3 model, although we conjecture that they maximise it.
IV.2 Coherent transport
Figure 3: Coherence element $|\rho_{C}^{1,3}|$ as a function of the system-environment coupling $\eta$ for different values of the effective temperature $\Omega$ when $J_{chain}=10$, $J_{res}=1$.
In Fig. 3 we plot $|\rho_{C}^{1,3}|$ after two time steps as a function of $\eta$ for different values of $\Omega$. The value of $|\rho_{C}^{1,3}|$ for $\eta=0$ represents the coherence between the first and last qubits in the absence of system-environment interaction. This value is obviously unaffected by the value of $\Omega$.
Moving on to values of $\eta<\frac{\pi}{4}$, which we shall call the “weak coupling” regime, we observe a decrease in $|\rho_{C}^{1,3}|$, meaning that system-environment interactions lead to less efficient coherent transport. Interestingly, for the same $\eta$, higher values of $\Omega$ lead to higher coherence, meaning that a Markovian environment is preferable to a non-Markovian one.
Approaching higher values of $\eta$, the coherence decreases monotonically for $\Omega=1$ while for $\Omega\neq 1$ it reaches a minimum and then increases again. The position of the minima depends on the value of $\Omega$: the more Markovian the environment, the stronger the system-environment coupling must be to minimise the coherence. In the “strong coupling” regime, we observe a considerable coherence revival, with lower values of $\Omega$ associated with higher rebounds.
So far, we have considered how the coherence is affected by $\eta$ and $\Omega$ after two iterations of the protocol. To check whether these results are consistent with the behaviour of the system over longer time scales, in Fig. 4 we plot $|\rho_{C}^{1,3}|$ as a function of the number of iterations of the protocol $n$ for different values of $\Omega$. We consider two cases: a) $\eta=0.4$ (weak coupling regime), and b) $\eta=1.2$ (strong coupling regime).
@@ IV.2 Coherent transport
Figure 4: Coherence element $|\rho_{C}^{1,3}|$ over multiple iterations of the protocol for different values of $\Omega$. The case $\Omega=0$ represents the unrealistic scenario of a fully unitary evolution. In this case, the coherence oscillates wildly, so the vertical axis has been truncated to improve the visibility of the graphs.
Starting with Fig. 4a, we observe that the first peak of $|\rho_{C}^{1,3}|$ is consistent with the results of Fig. 3: a Markovian environment results in greater coherence. Excluding the oscillatory behaviour of the ideal case $\Omega=0$ (no information loss), at later times the coherence is suppressed, with higher values of $\Omega$ generally associated with longer coherence times. This reinforces the interesting observation that a Markovian environment better assists coherent transport in the weak coupling regime.
@@
Consistently with Fig. 3, the situation is flipped in the strong coupling regime. In Fig. 4b we observe that the coherence is immediately suppressed for higher values of $\Omega$ while it is more persistent for highly non-Markovian environments.

### v3

@@ Abstract
We propose a collision model to investigate the information dynamics of a system coupled to an environment with varying degrees of non-Markovianity. We control the degree of non-Markovianity by applying a depolarising channel to a fixed and rigid reservoir of qubits. We characterise the effect of the depolarising channel and apply the model to study the coherent transport of information on a chain of three interacting qubits. We show how the system-environment coupling probability and the degree of non-Markovianity affect the process. Interestingly, in some cases a Markovian environment is preferable to reduce information loss and enhance the coherent transport.
@@ I Introduction
In this paper, we take an information-theoretic approach to investigate coherent transport in a simplified scenario. We propose a collision model [27, 28] to study the non-Markovian information dynamics of a chain of qubits interacting with the environment. Differently from other models in the literature [29, 30, 31, 32, 33, 34, 35, 36], the system interacts locally with the same reservoir qubits and we control the degree of non-Markovianity of the environment by applying a depolarising channel to the reservoir. By changing the intensity of the depolarising effect, we can control the amount of information irretrievably lost by the environment. We quantify this effect using a measure of non-Markovianity. We also study how non-Markovianity affects the rate of information loss and find a regime where Markovianity surprisingly reduces it.
We then apply the collision model to study the coherent transport of information, represented as an excitation, along a chain of qubits coupled to a fixed and rigid reservoir of interacting qubits. By changing the degree of non-Markovianity of the reservoir, we can model different types of environments. For example, a Markovian environment could represent a thermal bath surrounding the chain, while a non-Markovian one could model other degrees of freedom of the chain that are coupled to the ones responsible for the coherent transport. Our model is designed to study the dynamics of quantum information, not the spatial trajectory of a specific excitation, focusing specifically on the coherence of information transport along the chain.
We show that the transfer of information to the environment plays a crucial role in the coherence of the transport. The process displays different behaviours depending on the probability of the system-environment interaction. When the system-environment coupling is strong, non-Markovianity dramatically improves the coherence of the transport. Interestingly, when the coupling is weak, there is a regime where a Markovian environment is preferable to enhance the coherent transport.
Our results provide some information-theoretic insights into how non-Markovian environments affect coherent information transport on a chain. Moreover, our model demonstrates how to introduce a control on the degree of non-Markovianity in collision models with a fixed environment and might be applied in other fields where transport phenomena are relevant.
@@ II The model
We model the environment as a reservoir of $N$ interacting “reservoir qubits” $Q_{m}$, $m=1,...,N$. To allow the information to flow in the environment as well, their interaction is described by a Heisenberg Hamiltonian:
@@
with coupling constant $J^{m}_{res}\in\mathbb{R}\ \forall m=1,...,N$ and $\hat{\mathcal{X}}_{m}$, $\hat{\mathcal{Y}}_{m}$, $\hat{\mathcal{Z}}_{m}$ the Pauli operators of the m-th reservoir qubit. Since we are interested in how the information dynamics of the system is affected by a loss of information, we initialize the reservoir qubits in the maximally mixed state
@@
where $\mathbb{I}$ is the identity operator, $\hat{\mathcal{S}}_{m}$ the swap operator between the m-th system and reservoir qubit, and $\eta\in[0,\pi/2]$ a parameter that regulates the probability of the interaction. Since $\hat{P}_{m}(\eta)$ can be written as
@@
all the interactions in the model up to now are fundamentally of the same type. However, we prefer to maintain a continuous description in time for the chain and reservoir interactions to avoid introducing temporal order into the pairwise interactions between their qubits.
Since all the interactions considered so far are unitary and the same reservoir qubits keep interacting with the chain, the model is intrinsically non-Markovian. Unlike other non-Markovian collision models where the reservoir qubits that have interacted with the system are discarded at some point [32, 33, 34, 35, 36], we propose to control the degree of non-Markovianity of the model by applying a quantum depolarising channel to the fixed and rigid reservoir qubits after each collision. If $\rho_{CR}$ is the state of chain+reservoir, the depolarising channel for a single reservoir qubit can be expressed as:
@@
Figure 1: Four steps of the protocol with three chain qubits and three reservoir qubits. a) exchange step: each system qubit interacts with its nearest reservoir qubit via a partial swap; b) depolarising step: the depolarising channel is applied to the reservoir qubits; c) transfer via reservoir: unitary evolution of the reservoir qubits with $\hat{H}_{res}$; d) transfer via chain: unitary evolution of the system qubits with $\hat{H}_{chain}$.
@@
    The system qubits interact with the nearest reservoir qubits via a partial swap (Eq. (4)) with probability $\eta$ (Fig. 1a). We will call this step the “exchange step”.
@@
    The depolarising channel is applied to the reservoir qubits (Fig. 1b). This step controls the degree of non-Markovianity of the environment. We will call this step the “depolarising step”.
@@
Non-Markovianity.
We expect the parameter $\Omega$ to control the amount of information the environment loses. When $\Omega=0$, $\Delta_{0}^{m}(\rho_{CR})$ leaves $\rho_{CR}$ unchanged so that the global evolution of chain+environment is unitary. When $\Omega=1$, instead, $\Delta_{1}^{m}(\rho_{CR})$ re-initialises the $m$-th reservoir qubit to the maximally mixed state and, if applied to all the reservoir qubits, makes the environment Markovian. To quantify the effect of $\Omega$ on the degree of non-Markovianity, we use the measure proposed in [38]. Given two initial states of the system $\rho_{1}(0)$ and $\rho_{2}(0)$, the degree of non-Markovianity under the action of a dynamical map is measured as [38]:
@@
$\mathcal{N}=\underset{\{\rho_{1,0},\rho_{2,0}\}}{\text{max}}\sum_{{k}\in\Sigma_{+}}\left(D\left(\rho_{1,{k}},\rho_{2,{k}}\right)-D\left(\rho_{1,{k}-1},\rho_{2,{k}-1}\right)\right),$ (9)
where the summation is taken on all values of ${k}$ such that $D\left(\rho_{1,{k}},\rho_{2,{k}}\right)-D\left(\rho_{1,{k}-1},\rho_{2,{k}-1}\right)>0$, $\rho_{1,{k}}$ is the state of the system at the end of the $k$-th step obtained by tracing out the environment, and the maximisation is performed over all pairs of initial states $\{\rho_{1,0},\rho_{2,0}\}$.
@@
$\rho_{{l},0}=\frac{1}{2}\begin{pmatrix}1+r_{{l}}\,\text{cos}\left(\theta_{{l}}\right)&r_{{l}}\,\text{sin}\left(\theta_{{l}}\right)e^{-i\Phi_{{l}}}\\ r_{{l}}\,\text{sin}\left(\theta_{{l}}\right)e^{i\Phi_{{l}}}&1-r_{{l}}\,\text{cos}\left(\theta_{{l}}\right)&\end{pmatrix},$ (10)
@@
with $r_{{l}}\in\left[0,1\right]$, $\theta_{{l}}\in\left[0,\pi\right]$, and $\phi_{{l}}\in\left[0,2\pi\right]$, for ${{l}}=1,2$. The maximisation procedure becomes too complicated with a higher number of qubits. In this case, we use some of the insights on the states maximising Eq. (9) drawn from the 1x1 case to choose a suitable pair of initial states.
Figure 2: Degree of non-Markovianity $\mathcal{N}$ as a function of $\Omega$ for two different models. $\mathcal{N}$ is calculated over 30 time steps with $\eta=\pi/2$. The blue circles correspond to values of $\mathcal{N}$ calculated with 1000 randomly generated pairs of initial system states in the 1x1 model. The red circles are the values of $\mathcal{N}$ obtained with the initial states |$\ket{0}$ and |$\ket{1}$. The green squares correspond to the values of $\mathcal{N}$ for the 3x3 model with initial states |$\ket{000}$ and |$\ket{111}$. Note that the values of $\mathcal{N}$ when $\Omega\approx 0$ are bounded by the finite number of time steps of our simulations rather than by the state of the system reaching a steady state. Increasing the number of time steps would increase these values.
In Fig. 2, we plot the degree of non-Markovianity $\mathcal{N}$ induced by our protocol in two models with different numbers of qubits. The circles refer to a model with one system qubit and one reservoir qubit (1x1), and the squares to the model with three system qubits and three reservoir qubits (3x3). We have calculated $\mathcal{N}$ over 30 temporal steps with $\eta=\pi/2$ for different values of $\Omega$. The blue circles correspond to values of $\mathcal{N}$ for 1000 randomly generated pairs of initial states of the system in the 1x1 model. The red circles correspond to the initial states |$\ket{0}$ and |$\ket{1}$. The graph provides strong numerical evidence that |$\ket{0}$ and |$\ket{1}$ maximise $\mathcal{N}$ in Eq. (9). During the course of the simulations, we noticed that any two antipodal states on the Bloch sphere would maximise $\mathcal{N}$. This observation, and the fact that the 3x3 model consists of identical copies of the 1x1 model, lead us to consider |$\ket{000}$ and |$\ket{111}$ as good candidates to maximise $\mathcal{N}$ in the 3x3 model. The green squares correspond to values of $\mathcal{N}$ for this pair of initial states and show a behaviour consistent with the 1x1 case. Due to the computational demand of simulating the 3x3 model with many different initial states, we consider only this pair of initial states. More extensive simulations are required to confirm that these initial states maximise $\mathcal{N}$ as the considerations made for the 1x1 model might not apply to the 3x3 model. Therefore, the green squares can only be interpreted as a lower bound on the non-Markovianity of the 3x3 model, although we conjecture that they maximise it.
Information loss.
Due to the interaction with the environment, the system qubits will lose information in time. We expect the information loss rate to be related to the degree of non-Markovianity. To quantify this effect, we study the behaviour of the von Neumann entropy of a system qubit in a 1x1 model. We initialise the qubit in the state |$\ket{1}\bra{1}$ |, corresponding to a zero-entropy state (maximal information). In Fig. 3a, we plot the von Neumann entropy $S(\rho_{C}(k))$ of the system qubit after the $k$-th step of the protocol for $\eta=0.4$ and different values of $\Omega$. We observe that the entropy tends to 1 (no information) unless $\Omega=0$. If $\Omega=0$, the system will lose some information to the environmental qubit (which is initialised in the maximally mixed state and thus has no initial information) in a periodic process, but the information will never be irretrievably lost due to the overall unitary dynamics of system+environment. For any value of $\Omega>0$, the qubit will eventually lose all its information. Moreover, we note that the entropy for higher values of $\Omega$ increases more slowly than for lower values of $\Omega$, meaning that information loss is slower for lower degrees of non-Markovianity. Surprisingly, Markovianity protects the qubit from information loss, although it is associated with more information loss from the environment.
To study how this effect depends on $\eta$, we consider the time-average entropy of the qubit, defined as:
$\bar{S}(\eta,\Omega)=\frac{1}{K}\sum_{k=1}^{K}S(\rho_{C}(k)),$ (11)
where the average is taken over the first $K$ iterations of the protocol. We plot this quantity in Fig. 3b for $K=30$ and different values of $\Omega$. We exclude $\Omega=0$ as, in this case, the information is never entirely lost. Fig. 3b shows that, for $0.25\leq\Omega\leq 1$, lower degrees of non-Markovianity protect the qubit from information loss (lower values of $\bar{S}$) if system-environment interactions are weak (lower values of $\eta$). There is an inversion in this behaviour for higher values of $\eta$, although $\bar{S}$ is higher due to the qubit losing information to the environment more quickly in this regime. This protecting mechanism of Markovianity for small $\eta$ is surprising given that Markovianity is associated with complete information loss from the environment. We might expect this effect to favour the coherent transport of information on a chain of qubits. This will be the focus of the next Section.
Figure 3: (a) Von Neumann entropy of the systems qubit $S(\rho_{C}(k))$ over multiple iterations of the protocol for $\eta=0.4$ and different values of $\Omega$. (b) Time-averaged entropy $\bar{S}$ as a function of $\eta$ for different values of $\Omega$. The average is taken over 30 steps.
III Application to Coherent transport
@@
We apply the model described above to study the coherence of the transport of information along the chain. We represent initial information as a localized excitation
@@
in the Hilbert space $\mathscr{H}_{S_{1}}\otimes\dots\otimes\mathscr{H}_{S_{M}}$. Applying the protocol described above, the excitation spreads through the chain and the environment. After repeating the protocol $k$ times, we trace out the environment and extract the coherence between the first and last qubit of the chain $\rho_{C}^{1,N}(k)$, that is, the coefficient of the term |$\ket{100\dots 0}\bra{00\dots 01}$ | of the density matrix. We take $|\rho_{C}^{1,N}(k)|$ as an indicator of the coherence of the transport across the chain. We choose to study this parameter as the goal of this research is to show that the excitation transport along the chain occurs coherently, a feature that other entries of $\rho_{C}$ such as the population $\rho_{C}^{N,N}$ would not capture.
@@
For simplicity, we consider a model with three chain qubits. We implemented the protocol as a quantum circuit in Qiskit [39] as follows. The exchange phase consists of a sequence of local unitaries $\hat{P}(\eta)$ as in Eq. (4) for each of the system qubits $S_{i}$; the depolarising step is implemented as a depolarising noise model on the identity gate, which applies the quantum channel of Eq. (6); the transfer phases are implemented as unitaries on system and reservoir with a time step $\Delta t=0.01$. We choose $J^{n}_{chain}=J_{chain}=10,\,\forall n$ and $J^{m}_{res}=J_{res}=1,\,\forall m$. This choice represents a scenario where the chain is the most efficient medium for the propagation of the excitation.
We first apply the protocol 30 times and study how $|\rho_{C}^{1,3}(k)|$ varies for different values of $\Omega$ when $\eta=0.4$ and $\eta=1.2$. We select these values of $\eta$ as representative of the chain-environment “weak” and “strong” coupling regimes, respectively. This is because $\eta$ represents the probability of swapping the chain and reservoir qubits during the interaction described by Eq. (4): when $\eta<\frac{\pi}{4}\approx 0.8$, the probability is lower than $50\%$; when $\eta>\frac{\pi}{4}\approx 0.8$ the probability is higher than $50\%$. After studying the temporal behaviour of $|\rho_{C}^{1,3}(k)|$ at fixed $\eta$, we investigate the effect of $\eta$ on the coherence of the transport. We do so by examining the maximum coherence $\max_{k}|\rho_{C}^{1,3}(k)|$ as a function of $\eta$ for different values $\Omega$. We then restrict again to the cases $\eta=0.4$ and $\eta=1.2$ to study how the ratio $J_{res}/J_{chain}$ affects $\max_{k}|\rho_{C}^{1,3}(k)|$. Finally, we examine the temporal behaviour of the populations $|\rho_{C}^{1,1}|$ and $|\rho_{C}^{3,3}|$ in the same conditions, that is, the coefficients of the terms |$\ket{100}\bra{100}$ | and |$\ket{001}\bra{001}$ |, respectively.
@@
Figure 4: Coherence element $|\rho_{C}^{1,3}(k)|$ over multiple iterations of the protocol for different values of $\Omega$. The case $\Omega=0$ represents the unrealistic scenario of a fully unitary evolution. In this case, the coherence oscillates wildly, so the vertical axis has been truncated to improve the visibility of the graphs.
In Fig. 4 we plot $|\rho_{C}^{1,3}(k)|$ as a function of the number of iterations of the protocol $k$ for different values of $\Omega$. We consider two cases: a) $\eta=0.4$ (weak coupling regime) in Fig. 4a, and b) $\eta=1.2$ (strong coupling regime) in Fig. 4b.
Before analysing the two cases separately, we notice that in both scenarios $|\rho_{C}^{1,3}(k)|$ is not affected by $\Omega$ for $k=1$, i.e., considering a protocol with a single iteration. This is because each protocol iteration contains a single interaction between the chain and the environment (described by $\hat{P}_{m}(\eta)$ in Eq. (4)), so there is no further exchange of information after the application of the depolarising channel. This makes the degree of non-Markovianity of the environment irrelevant to the coherent transport on the chain.
In both Fig. 4a and Fig. 4b, $\Omega=0$ corresponds to a fully unitary evolution of the global system chain+environment. In this case, no information is lost, and no steady state can be reached. Such a scenario is unrealistic in most ordinary physical systems, where some information is irreparably lost over time. For this reason, we will consider only $\Omega\geq 0.25$ in the following discussion. In the cases examined in Figs. 4a-4b we see that $|\rho_{C}^{1,3}(k)|$ is quickly damped and reaches zero within 30 iterations of the protocol. We expect this to happen for any $\Omega\neq 0$ due to the depolarising channel losing information at each protocol iteration, although the damping will take longer for values of $\Omega$ close to zero.
@@
Focusing now on Fig. 4a, we observe that, in the weak coupling regime $\eta=0.4$, higher values of $\Omega$ result in greater coherence. This is analogous to what we observed in Sec. II and it is interesting as one may expect more information loss to be associated with decreasing amounts of coherence in the system. Instead, we observe that, for $0.25\leq\Omega\leq 1$, a Markovian environment better assists coherent transport. The situation is inverted in the strong coupling regime. In Fig. 4b, we observe that the coherence is immediately suppressed for higher values of $\Omega$ while it is higher and more persistent for highly non-Markovian environments.
Figure 5: Maximum coherence $\max_{k}|\rho_{C}^{1,3}(k)|$ over $k=30$ protocol iteration as a function of the system-environment coupling $\eta$, for different values of $\Omega$ when $J_{chain}=10$, $J_{res}=1$. The case $\Omega=1$ tends to zero when $\eta\rightarrow\frac{\pi}{2}$, so the y-axis has been truncated to improve the visibility of the results.
So far, we have considered how $|\rho_{C}^{1,3}(k)|$ is affected by $\Omega$ for fixed values of the interaction probability $\eta$, one in the weak coupling regime and the other in the strong coupling regime. To check how these results depend on the value of $\eta$, in Fig. 5 we plot the maximum of the coherence $\max_{k}|\rho_{C}^{1,3}(k)|$ computed over 30 protocol iterations as a function of $\eta$ for different values of $\Omega\geq 0.25$. The results are consistent with what was shown above: the coherence slightly increases with the Markovianity of the environment in the weak coupling regime, while a non-Markovian environment is preferred in the strong coupling regime. The transition between these two regimes happens for intermediate values of the coupling probability that depend on $\Omega$. For $\eta=0$, since there is no coupling between the chain and the environment, $\max_{k}|\rho_{C}^{1,3}(k)|$ does not depend on $\Omega$. For $\eta$ close to $\frac{\pi}{2}$, i.e., with an interaction probability close to 1, the coherence goes to zero for $\Omega=1$. In this case, the coherence is immediately suppressed during the first protocol iteration because the excitation is entirely transferred to the reservoir and erased in the first two steps of the protocol (exchange and depolarising). For $\Omega\neq 1$, $\text{max}_{k}|\rho_{C}^{1,3}|$ is always nonzero and, after reaching a minimum for some intermediate values of $\eta$, it increases when $\eta$ tends to $\frac{\pi}{2}$, with higher values of coherence associated with lower values of $\Omega$.
(a) $\eta=0.4$
(b) $\eta=1.2$
Figure 6: Maximum coherence $\max_{k}|\rho_{C}^{1,3}(k)|$ over $k=30$ protocol iteration as a function of the ratio between the reservoir qubits and the chain qubits couplings $J_{res}/J_{chain}$, for $J_{chain}=10$ and for different values of the degree of non-Markovianity $\Omega$.
Our results so far are based on simulations with $J_{chain}=10$ and $J_{res}=1$. As we mentioned above, this corresponds to a scenario where the chain is the most efficient medium for the propagation of the excitation. While this situation is the most physically realistic, in Fig. 6a we explore how our results depend on the ratio $J_{res}/J_{chain}$. Specifically, we plot $\max_{k}|\rho_{C}^{1,3}(k)|$ as a function of $J_{res}/J_{chain}$ in the weak ($\eta=0.4$) and strong ($\eta=1.2$) coupling regimes for $J_{chain}=10$ and different values of $\Omega\geq 0.25$. As expected, $J_{res}/J_{chain}$ does not affect $\max_{k}|\rho_{C}^{1,3}(k)|$ when $\Omega=1$ since no transfer of information along the environment is possible when all the information stored in the environment is completely lost at each protocol iteration.
Excluding the case $\Omega=1$, in Fig. 6a we observe a low sensitivity of $\max_{k}|\rho_{C}^{1,3}(k)|$ to changes in $J_{res}/J_{chain}$, with higher values of $\Omega$ initially associated to higher levels of coherence, consistently with Figs. 4-5. The low sensitivity to $J_{res}/J_{chain}$ indicates that the chain is the main responsible for the coherent transport of the excitation. When $\Omega=0.25$ we observe a clear-cut change in behaviour for $J_{res}/J_{chain}=0.5$. After this point, $\max_{k}|\rho_{C}^{1,3}(k)|$ increases more rapidly with increasing values of $J_{res}/J_{chain}$, suggesting a more prominent role of the intra-environment interactions in the coherent transport.
The situation is different in Fig. 6b, where we observe that, apart from the case $\Omega=1$ discussed above, $\max_{k}|\rho_{C}^{1,3}(k)|$ consistently increases with $J_{res}/J_{chain}$, thus pointing to the active role of the environment in the coherent transport of the excitation. This happens because the environment is more likely to interact with the chain in the strong coupling regime and is thus more likely to assist in the propagation of the excitation. Accordingly, higher degrees of non-Markovianity result in higher values of coherence since the information in the environment is less likely to be lost.